OpenStudy (jazzyluvsu123):

Ok i know how to do Order Of Operations but I cant seem to get these 3 1. (11+7) /9+3-2 2. (14-6) exponent 2 -8 +4x3 3. (12+8)/5---- answer divided by 2

OpenStudy (mertsj):

1. 18/9+3-2 2+3-2=3

OpenStudy (mertsj):

2. 8^2-8+4x3 64-8+12 56+12=68

OpenStudy (mertsj):

3. 20/5divided by 2 4divided 2 = 2

OpenStudy (jazzyluvsu123):

thanks im gonna give u medal

OpenStudy (mertsj):

ty